. 2 L D5RL at 150 mL/hr. Drop factor: 15 gtt/mL

what is the IV Flow rate in gtt/min
Answer #1

2000ml solution at 150ml/hr flow rate and drop factor 15gtt/ml

what is flow rate in gtt/min ?

formula:

mL/hr X drop factor = gtt/min.

150 ml/hr X15gtt/min= 2250/60=37.5 gtt/min will be IV flow rate in gtt/min